OpenStudy (anonymous):

Gasoline is pouring into a cylindrical tank of radius 3 feet. When the depth of the gasoline is 6 feet, the depth is increasing at 0.2 ft/sec. How fast is the volume of gasoline changing at that instant? Round your answer to three decimal places.

OpenStudy (anonymous):

Okay, you can find the volume first since you have h and r , so : \[V = \pi r^2 h\] \[= ( 3)^2 (6) \pi\] = 169.65 ft^3

OpenStudy (anonymous):

sstarica, isn't this the same thing we were tackling

OpenStudy (anonymous):

now find V' which is :\[V' = 2\pi rr'h'\]

OpenStudy (anonymous):

not really @iam , somehow similar :)

OpenStudy (anonymous):

now we have: h' = 0.2 ft/sec we need to find r'

OpenStudy (anonymous):

do i just plug in

OpenStudy (anonymous):

yes :) but we need to find r' now

OpenStudy (anonymous):

\[V' = 2 \pi rr' h'\] we have h' , but we need to find r' to calculate V' = rate of volume increase

OpenStudy (anonymous):

how do i do that? :(

OpenStudy (anonymous):

hold on :)

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

you need to find a relationship between r and h so r/h = r'/h'

OpenStudy (anonymous):

thats where im having problem for all the problems

OpenStudy (anonymous):

cross mulitply and you'll get: r' = rh'/r

OpenStudy (anonymous):

sorry! wait you'll get: r' = rh'/h :)

OpenStudy (anonymous):

are you with me soo far?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

now plug : r =3 ft h = 6ft h' = 0.2ft/sec r' = (3)(0.2)/6 = 0.1 ft/sec

OpenStudy (anonymous):

we have found r' now, let's go back to V'\[V' = 2 \pi rr'h'\] so, V' = 2(3)(0.1)(0.2) pi = 0.377 ft^3/sec

OpenStudy (anonymous):

Correct me if I'm wrong please

OpenStudy (anonymous):

OMG this makes so much sense now

OpenStudy (anonymous):

but the website marked it wrong :(

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

I guess there's a mistake, but that's the idea, maybe in calculation?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

did you get the idea?

OpenStudy (anonymous):

one qquestion

OpenStudy (anonymous):

sure :)

OpenStudy (anonymous):

where did u get \[V \prime = 2\pi r r \prime h \prime \]

OpenStudy (anonymous):

you derive with respect to r and h :)

OpenStudy (anonymous):

you want to know how is the volume increasing when both the radius and height is changing, that's the story of the related rated problem

OpenStudy (anonymous):

oooooh...im sorry can u show me one more problem

OpenStudy (anonymous):

hmmmm, let's say we have a circle of radius = 2 m, and the radius is decreasing at 0.5 m/sec. How much will the circle's area increase ?

OpenStudy (anonymous):

I'll give you advise, whenever you see ( m/sec) or anything with respect to time it's something' which in this case is the r' = 0.5m/sec

OpenStudy (anonymous):

so what is the given?

OpenStudy (anonymous):

we have : - R = 2 m - R' = 0.5 m/sec and we know the formula of the Area of the circle which is: A = pi r^2, right?

OpenStudy (anonymous):

uhhuh

OpenStudy (anonymous):

Now , you want to know how much is the area increasing with respect to R, since R is increasing :) so you find the derivative of it which will be : A' = 2 pi rr'

OpenStudy (anonymous):

so when exactly do i start pluging in numbers?

OpenStudy (anonymous):

right now, after you find the derived formula, start pluggingthe numbers :)

OpenStudy (anonymous):

okay ...these are pretty simple compared to the ones i have on hw :'(

OpenStudy (anonymous):

Like this problem, they want to know how much is the area increasing , we got the formula so the answer is: A' = 2(2)(0.5) pi = 2 pi = 2.68 m^2/sec :) So the area is increasing at 2.68 m^2/sec

OpenStudy (anonymous):

Lol, when you understand the small idea, you'll be able to figure out the bigger idea :)

OpenStudy (anonymous):

did you understand my problem?

OpenStudy (anonymous):

yes but how would i approach towards a hemisphere problem

OpenStudy (anonymous):

think about it, write down the given on one side, and the RTF(required to find) on the other side, then SKETCH! To test your understanding , you've got to sketch what they give you to be fully understand what they want :)

OpenStudy (anonymous):

to fully*

OpenStudy (anonymous):

1) write the given and what you need to find 2) sketch 3) write down the general formula that you're going to use :)

OpenStudy (anonymous):

^_^

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

good luck ! You're a smart student, you can figure it out :)

OpenStudy (anonymous):

you got the small idea, you'll be able to solve any bigger ideas now . They are not even big, they are small ideas that got stretched ^_^, don't let the problem boss u, YOU ARE THE BOSS!

OpenStudy (anonymous):

lol..thx

OpenStudy (anonymous):

np :) Give it a try once more. ^_^